In a partially destroyed laboratory, record of an analysis of correlation of data,
only the following results are legible:
Variance of X = 9
Regression equations are
(i) 8x −10y + 66 = 0
(ii) 0 40x −18y − 214 =
Find out the following missing results.
(i) The means of X and Y
(ii) The coefficient of correlation between x and y
(iii) The standard deviation of Y
1
Expert's answer
2019-03-14T15:26:29-0400

Since two regression lines always intersect at a point representing mean values of the values xmean and ymean


"8x-10y+66=0""40x-18y-214=0"

"8x-10y=-66"

"32y=544 => ymean=17"

"xmean=(10*17-66)\/8=13"

"xmean=13, ymean=17"

(ii) To find the given regression equations in such a way that the coefficient of dependent variable is less than one at least in one equation.


"8x-10y=-66 => y=66\/10+8x\/10"

byx = 0.8


"40x-18y=214 => x=214\/40+18y\/40"

bxy = 0.45


"r = sqrt ( bxy * byx )"

"r = sqrt(0.45 * 0.8)=0.6"

Coefficient of correlation r between x and y is 0.6.

(iii) To determine the standard deviation of y , consider the formula:


"\sigma y = (\sigma x * byx)/r"

"\u03c3y= 0.8 * sqrt(9) \/ 0.6 = 4"